OpenStudy (anonymous):

Which relation is a direct variation that contains the ordered pair (2, 7)?

OpenStudy (anonymous):

is this a multiple choice question? can you provide the choices?

OpenStudy (anonymous):

Y=4x-1 y=2/7x y=7/2x y=7/x

OpenStudy (anonymous):

the third one

OpenStudy (anonymous):

The graph of a function must be linear if it has what characteristic?

OpenStudy (anonymous):

the x is raised to the power 1 and not more!

OpenStudy (anonymous):

A salesperson earns $200 a week plus a 4% commission on her sales. Which equation models the relationship between her sales in a given week, s, and her weekly earnings, e?

OpenStudy (anonymous):

200s+4/100 4s+200 e=4/100s+200 200s+4